OpenStudy (anonymous):

At a video rental store, Elyssa has only enough money to rent three videos. She has chosen four comedies, six dramas, and one mystery movie to consider. Part A) How many different selections of three videos may she rent from the movies she has chosen? Part B) How many selections of three videos will consist of one comedy and two dramas? Part C) The probability that a selection of three videos will consist of one of each type of video is... Part D) Elyssa decides to rent one comedy, one drama, nd one mystery movie. She may view the videos in ___ different orders.

OpenStudy (anonymous):

\(\dbinom{11}{3}\) for the first one, which doesn't really answer the question, it just restates it since this means literally "how many ways to choose 3 from a set of 11" now you have a choice you can learn how to compute this (it is not hard) or you can use a calculator (almost all have this feature) or while you are on the computer you can do this http://www.wolframalpha.com/input/?i=11+choose+3

OpenStudy (anonymous):

one comedy out of 4, 4 ways two dramas out of 6 \(\binom{6}{2}=\frac{6\times 5}{2}=15\) and \[4\times 15=60\]